OpenStudy (anonymous):

s(t)=68-20log(t+1),t>0

OpenStudy (anonymous):

what is the actual question?

OpenStudy (dumbcow):

If you are trying to graph this, notice if t=0 then s(0) = 68 As t increases, log(t+1) increases resulting in s(t) to decrease indefinitely to find x-intercept, set s(t) = 0, solve for t 20log(t+1) = 68 log(t+1) = 3.4 (I will assume this is natural log) t+1 = e^3.4 t = 28.96